OpenStudy (anonymous):

-3x < -6 or x + 5 < -2

OpenStudy (anonymous):

-3x < -6 or x + 5 \[\le\] -2***************************

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

yes im so confused. and frustrated

OpenStudy (mysesshou):

i dont remember what the OR means for these. \[-3x < -6\] or \[x + 5 \le -2 \]

OpenStudy (anonymous):

im not sure either i guess solve both?

OpenStudy (anonymous):

In algebra, disjunction refers to the combining of math sentences with the word "OR." Some refer to this as a union of two or more sentences.

OpenStudy (mysesshou):

I guess it is more of the same. So for the first one, what would you do first?

OpenStudy (anonymous):

divide -6 by -3

OpenStudy (mysesshou):

Yes, and also flip the symbol, since dividing by negative

OpenStudy (mysesshou):

which is ?

OpenStudy (anonymous):

x>2

OpenStudy (mysesshou):

yes. ok, next one.

OpenStudy (mysesshou):

to isolate the x, you should do what?

OpenStudy (anonymous):

x\[\le\]-7?

OpenStudy (mysesshou):

Yes, \[x \le -7\] by subtracting the five each side

OpenStudy (anonymous):

then graph

OpenStudy (mysesshou):

So, \[x \le -7\] and x>2 Let's flip the one with the -7 in it to give \[-7 \ge x\]

OpenStudy (mysesshou):

hang on. let me check that

OpenStudy (anonymous):

ok

OpenStudy (mysesshou):

Ok, this one doesn't combine. x≤−7 and x>2

OpenStudy (anonymous):

how would i graph this?

OpenStudy (mysesshou):

So, filled in circle at -7, and open circle at 2. Then since x is smaller than -7, draw arrow to the left, for the values less than -7. also draw arrow to the right from the 2, showing values above 2

OpenStudy (anonymous):

thank you :))

OpenStudy (mysesshou):

welcome :)